# August 2025 IPMU25-0038

### v1

@@ 4.1 Einstein frame
$\varepsilon^{(E)}={}-{{1}\over{H_{E}^{2}}}{{dH_{E}}\over{dt_{E}}}$ (35)
@@
$\eta^{(E)}=\varepsilon^{(E)}-{{1}\over{2\,\varepsilon^{(E)}\,H_{E}}}{{d\varepsilon^{(E)}}\over{dt_{E}}}~.$ (36)
After the differentiation of Eq. (35) as
${{d\varepsilon^{(E)}}\over{dt_{E}}}=2H_{E}\,\left(\varepsilon^{(E)}\right)^{2}-{{1}\over{H^{2}_{E}}}{{d^{2}H_{E}}\over{dt_{E}^{2}}}~,$ (37)
@@
$\eta^{(E)}={}-{{1}\over{2\,H_{E}}}{{{{d^{2}H_{E}}\over{dt_{E}^{2}}}}\over{{{dH_{E}}\over{dt_{E}}}}}\,~.$ (38)
@@
${{d^{2}H_{E}}\over{dt_{E}^{2}}}={}-{{1}\over{M^{2}_{PL}}}{{d^{2}\phi}\over{dt_{E}^{2}}}\,{{d\phi}\over{dt_{E}}}\,$~, (39)
@@
$\eta^{(E)}={}-{{1}\over{H_{E}{{d\phi}\over{dt_{E}}}}}{{d^{2}\phi}\over{dt_{E}^{2}}}~.$ (40)
Next, substituting Eq. (40) into Eq. (34) yields the inflaton equation of motion in the form  [^4] Equation (41) in the case of the ultra-slow-roll (USR) regime with $V_{E,\phi}\longrightarrow 0$ implies $\eta^{(E)}\longrightarrow 3$. In the literature, the normalization of the SR parameter $\eta^{(E)}$ sometimes differs by the sign and the factor of 2.
$H_{E}{{d\phi}\over{dt_{E}}}\left(3-\eta^{(E)}\right)+V_{E,\phi}=0~.$ (41)
@@
in terms of e-folds $dN_{E}=-H_{E}dt=-dk/k$ or scale $k$. This is consistent with $N_{E}=-\ln(a_{E}/a_{E0})$, where $a_{E}(t)$ is the FLRW scale factor in the Einstein frame and the constant $a_{E_{0}}$ is the value of $a_{E}$ at the end of inflation. Hence, $N_{E}$ is a monotonically decreasing function during inflation and $N_{E}=0$ corresponds to the end of inflation. The Planck pivot scale $k_{*}$ of the horizon crossing is taken at $k_{*}=0.05~{\rm Mpc}^{-1}$, which corresponds to $N_{E*}=56\pm 8$ for Starobinsky inflation [38].
@@ 4.2 Jordan frame
$\varepsilon_{1}={}-{{\dot{H}}\over{H^{2}}}={{1}\over{2}}\,{{d\ln(H^{2})}\over{dN}}\,,\qquad\varepsilon_{n}={{\dot{\varepsilon}_{n-1}}\over{H{\varepsilon}_{n-1}}}={}-{{d\ln(\varepsilon_{n-1})}\over{dN}}~.$ (43)
We find useful to introduce the additional SR parameters in our setup, which are given by
$\zeta_{1}={{\dot{F}_{,R}}\over{HF_{,R}}}={{F_{,RR}\dot{R}}\over{HF_{,R}}}={}-{{d\ln(F_{,R})}\over{dN}}\,,\quad\zeta_{n}={{\dot{\zeta}_{n-1}}\over{H{\zeta}_{n-1}}}={}-{{d\ln(\zeta_{n-1})}\over{dN}}\,.$ (44)
Using Eq. (10), the first SR parameters can be expressed in terms of $H^{2}$ and $R$ as
$\varepsilon_{1}=2-{{R}\over{6H^{2}}},\quad\zeta_{1}={{\left(6H^{2}-R\right)F_{,R}+F}\over{6H^{2}F_{,RR}\left(12H^{2}-R\right)}}-{{R}\over{3H^{2}}}~.$ (45)
Equation (11) can be rewritten to the form
$Y={{M_{\mathrm{Pl}}H}\over{\sqrt{2F_{,R}\,}}}\left(1+{{1}\over{2}}\zeta_{1}\right)~,$ (46)
and Eq. (13) can be rewritten to
$3M_{\mathrm{Pl}}^{4}H^{2}\left(1+{{1}\over{2}}\zeta_{1}\right)^{2}=AF_{,R}{\dot{R}}^{2}+2F_{,R}V_{\mathrm{eff}}~.$ (47)
@@
${{dH_{E}}\over{dt_{E}}}={{M_{\mathrm{Pl}}\dot{Y}}\over{\sqrt{2\,F_{,R}}}}={{M_{\mathrm{Pl}}^{2}H^{2}}\over{8F_{,R}}}\left(2\zeta_{1}\zeta_{2}-(2\varepsilon_{1}+\zeta_{1})(2+\zeta_{1})\right)\,$, (48)
@@
$\begin{split}\varepsilon^{(E)}&={}-{{1}\over{H_{E}^{2}}}\,{{d{H}_{E}}\over{d{t}_{E}}}=\varepsilon_{1}+{{\zeta_{1}\left(1-\varepsilon_{1}\right)}\over{2+\zeta_{1}}}-{{2\zeta_{1}\zeta_{2}}\over{\left(2+\zeta_{1}\right)^{2}}}\\ &\approx\varepsilon_{1}+{{1}\over{2}}\zeta_{1}-{{\zeta_{1}}\over{4}}\left(\zeta_{1}+2\zeta_{2}+2\varepsilon_{1}\right),\end{split}$ (49)
@@
Using Eq. (19) in the following form:
$\zeta_{1}={}-2\varepsilon_{1}+\zeta_{1}\left(\zeta_{1}+\zeta_{2}-\varepsilon_{1}\right)~,$ (50)
@@
$\varepsilon^{(E)}\approx\varepsilon_{1}+{{1}\over{2}}\zeta_{1}-{{\zeta_{1}}\over{4}}\left(\zeta_{1}+2\zeta_{2}+2\varepsilon_{1}\right)={{1}\over{4}}\zeta_{1}\left(\zeta_{1}-4\varepsilon_{1}\right)\approx 3\varepsilon_{1}^{2}+{\cal O}\left(\varepsilon_{1}^{3}\right).$ (51)
Similarly, by using Eqs. (36) and (51) with
${{dN}\over{dN_{E}}}={{2}\over{2+\zeta_{1}}}\approx 1+\varepsilon_{1}\,$, (52)
we get
$\begin{split}\eta^{(E)}&=\varepsilon^{(E)}+{{1}\over{2}}\,{{d\ln(\varepsilon^{(E)})}\over{dN_{E}}}\\ &\approx{}-\varepsilon_{2}{{dN}\over{dN_{E}}}+3\varepsilon_{1}^{2}+{\cal O}\left(\varepsilon_{1}^{3}\right)={}-\varepsilon_{2}-\varepsilon_{1}\varepsilon_{2}+3\varepsilon_{1}^{2}+{\cal O}\left(\varepsilon_{1}^{3}\right)\,.\end{split}$ (53)
@@
$r\!$ $\approx$ $\!\!16\varepsilon^{(E)}\approx 4\left|\zeta_{1}\left(\zeta_{1}-4\varepsilon_{1}\right)\right|\approx 48\varepsilon_{1}^{2}\,$, (54)
$n_{s}\!$ $\approx$ $\!\!\!\!\!1-4\varepsilon^{(E)}+2\eta^{(E)}\!\approx\!\!1-2\varepsilon_{2}-6\varepsilon_{1}^{2}-2\varepsilon_{1}\varepsilon_{2}\!\approx\!\!1-2\varepsilon_{2}\,$, (55)
$\alpha_{s}\!$ $\approx$ $\!\!\!{}{}-2\varepsilon_{2}\varepsilon_{3},\qquad\alpha_{t}\approx 2~{{d\,\varepsilon^{(E)}}\over{dN_{E}}}\approx{}-12\varepsilon^{2}_{1}\varepsilon_{2}.$ (56)
@@
$A_{s}={{H_{E}^{2}}\over{8\pi^{2}M_{\mathrm{Pl}}^{2}\varepsilon^{(E)}}}\approx{{H^{2}}\over{48\pi^{2}F_{,R}\varepsilon^{2}_{1}}}\approx{{H^{2}}\over{\pi^{2}rF_{,R}}}\,.$ (57)
Of course, Eqs. (5457) make sense only in the SR approximation, because the higher-order terms with respect to $\varepsilon^{(E)}$ and $\eta^{(E)}$ are ignored.
@@ 4.3 USR regime
$A\ddot{\sigma}+{{A_{,\sigma}}\over{2}}{\dot{\sigma}}^{2}\ll V_{\mathrm{eff},\sigma}$ (58)
because the field derivative of the potential $V_{\mathrm{eff}}$ almost vanishes. Instead, Eq. (28) reads
$\ddot{\sigma}={}-3Y\sqrt{{{2F_{,R}}\over{M_{\mathrm{Pl}}^{2}}}}\dot{\sigma}-{{A_{,\sigma}}\over{2A}}{\dot{\sigma}}^{2}.$ (59)
The USR regime can be realized by engineering a nearly-inflection point in the inflaton potential below the inflation scale. It follows from Eq. (9) that at such (de Sitter-like) point one has
$H_{dS}^{2}={{F(R_{dS})}\over{6F_{,R}(R_{dS})}}\quad{\rm and,~hence,}\quad R_{dS}={{2F(R_{dS})}\over{F_{,R}(R_{dS})}}~.$ (60)
Under the condition (6), Eq. (60) is equivalent to
$V_{eff,\sigma}(\sigma_{dS})=0\quad{\rm and,~hence,}\quad F_{,\sigma}(\sigma_{dS})={{2F(\sigma_{dS})}\over{\sigma_{dS}}}\,.$ (61)
@@
$F_{,\sigma\sigma}(\sigma_{dS})={{2F(\sigma_{dS})}\over{\sigma_{dS}^{2}}}\,.$ (62)
@@
$V_{E}={{M^{2}_{\mathrm{Pl}}}\over{2}}R\exp\left(-\sqrt{{{2}\over{3}}}{{\phi}\over{M_{\mathrm{Pl}}}}\right)-F\exp\left(-2\sqrt{{{2}\over{3}}}{{\phi}\over{M_{\mathrm{Pl}}}}\right)~,$ (63)
$V_{E,\phi}=\sqrt{{{2}\over{3}}\,}{M_{\mathrm{Pl}}}^{-1}\left({F}\exp\left(-2\sqrt{{{2}\over{3}}}{{\phi}\over{M_{\mathrm{Pl}}}}\right)-V_{E}\right)~,$ (64)
@@
${V_{E,\phi\phi}}$ ${{}=\sqrt{{{2}\over{3}}}\left[{{M_{\mathrm{Pl}}}\over{2}}\exp\left(-\sqrt{{{2}\over{3}}}{{\phi}\over{M_{\mathrm{Pl}}}}\right){{dR}\over{d\phi}}\right.}$
${{}\left.{}-{{2F}\over{M^{2}_{\mathrm{Pl}}}}\sqrt{{{2}\over{3}}}\exp\left(-2\sqrt{{{2}\over{3}}}{{\phi}\over{M_{\mathrm{Pl}}}}\right)-{{V_{E,\phi}}\over{M_{\mathrm{Pl}}}}\right]~.}$
@@
$V_{E}\approx F\exp\left(-2\,\sqrt{{{2}\over{3}}}{{\phi}\over{M_{\mathrm{Pl}}}}\right)\approx{{F}\over{F_{,R}^{2}}}~.$ (66)
@@ 5 More on SR approximation in Jordan frame
Equation (13) in the SR approximation is just the Friedmann equation,
$3M_{\mathrm{Pl}}^{2}Y^{2}\approx V_{\mathrm{eff}}\,.$ (67)
After differentiating this equation with respect to the cosmic time in the Jordan frame and using Eq. (29), we obtain
$\dot{R}\approx{}-{{M_{\mathrm{Pl}}V_{\mathrm{eff},R}}\over{3YA\sqrt{2F_{,R}}}}\,~.$ (68)
@@
${R\,}^{\prime}\approx{{M_{\mathrm{Pl}}V_{\mathrm{eff},R}}\over{3YHA\sqrt{2F_{,R}}}}={{M_{\mathrm{Pl}}V_{\mathrm{eff},R}Y}\over{3Y^{2}HA\sqrt{2F_{,R}}}}\approx{{M_{\mathrm{Pl}}^{4}V_{\mathrm{eff},R}}\over{2F_{,R}AV_{\mathrm{eff}}}}\left(1-{{F_{,RR}}\over{2F_{,R}}}{R\,}^{\prime}\right)\,.$ (69)
After solving Eq. (69) for ${R\,}^{\prime}$, we get the SR evolution equation for the Ricci scalar $R$ in the Jordan frame as
${R\,}^{\prime}\approx{}-{{2F_{,R}\left(RF_{,R}-2F\right)}\over{F_{,RR}\left(2RF_{,R}-F\right)}}\,.$ (70)
@@
$\zeta_{1}(R)\approx{{2\left(RF_{,R}-2F\right)}\over{2RF_{,R}-F}}$ (71)
@@
$H^{2}(R)\approx{{\left(2RF_{,R}-F\right)^{2}}\over{54F_{R}\left(RF_{R}-F\right)}}~.$ (72)
Equation (45) also implies
$\varepsilon_{1}=2-{{R}\over{6H^{2}}}\approx 2-{{9F_{R}R\left(RF_{R}-F\right)}\over{\left(2RF_{,R}-F\right)^{2}}}\,~.$ (73)
Using Eqs. (71) and (73), we find in the SR approximation that
$\varepsilon_{1}(R)={}-{{1}\over{2}}\zeta_{1}(R)+{{1}\over{4}}\zeta_{1}^{2}(R).$ (74)
Then, by using Eqs. (70), (71) and (73), we get
$\varepsilon_{2}(R)={}-{{1}\over{\varepsilon_{1}}}{{d\varepsilon_{1}}\over{dN}}={}-{{{R\,}^{\prime}}\over{\varepsilon_{1}}}{{d\varepsilon_{1}}\over{dR}}\\ \approx{{18FF_{,R}\left(FF_{,RR}R-F_{,R}^{2}R+F_{,R}F\right)}\over{F_{,RR}\left(4F_{,R}^{3}R^{3}-3F^{2}F_{,R}R+F^{3}\right)}}~,$ (75)
$\zeta_{2}(R)\approx{{6F_{,R}\left(\left[FF_{,RR}-F_{,R}^{2}\right]R+F_{,R}F\right)}\over{F_{,RR}\left(2F_{,R}R-F\right)^{2}}}~,$ (76)
@@
$\begin{split}\zeta_{3}(R)&\approx{}-{{4\left(F_{,R}R-2\,F\right)}\over{F_{,RR}^{2}\left(F_{,RR}FR-F_{,R}^{2}R+F_{,R}F\right)\left(2F_{,R}R-F\right)^{2}}}\\ &\times\left(F_{,RR}\left[{{F}\over{2}}F_{,RR}^{2}R\left(2F_{,R}R+F\right)-{{F}\over{2}}F_{,R}F_{,RR}\left(2F_{,R}R-3F\right)\right.\right.\\ &\left.\left.{}+F_{,R}^{3}\left(F-F_{,R}R\right)\right]-{{1}\over{2}}F_{,RRR}F_{,R}^{2}\left(F_{,R}R-F\right)\left(2F_{,R}R-F\right)\right)\,.\end{split}$ (77)
Equation (74) in the same approximation implies
${{d\epsilon_{1}}\over{dN}}\approx-{{1}\over{2}}{{d\zeta_{1}}\over{dN}}+{{1}\over{4}}{{d\zeta_{1}^{2}}\over{dN}}\approx{{\zeta_{1}\zeta_{2}}\over{2}}-{{\zeta_{1}^{2}\zeta_{2}}\over{2}}\approx{{\zeta_{1}}\over{2}}\left(\zeta_{2}-\zeta_{1}\zeta_{2}\right)~.$ (78)
@@
$\epsilon_{2}\approx\zeta_{2}(1-\zeta_{1})\approx\zeta_{2}$ (79)
@@
$\epsilon_{3}\approx\zeta_{3}-{{\zeta_{1}\zeta_{2}}\over{1-\zeta_{1}}}\approx\zeta_{3}.$ (80)
We use the simplified relations $\varepsilon_{1}\approx{}-\zeta_{1}/2$, $\varepsilon_{2}\approx\zeta_{2}$, and $\varepsilon_{3}\approx\zeta_{3}$ to calculate the inflationary parameters by Eqs. (5456), and Eq. (73) to get the value of $R$ corresponding to the end of inflation. Substituting those results into the right-hand-side of Eq. (54) gives
$r\approx 48\varepsilon_{1}^{2}\approx 12\zeta_{1}^{2}\approx 48\left({{RF_{,R}-2F}\over{2RF_{,R}-F}}\right)^{2}$ (81)
@@
$\begin{split}n_{s}\approx 1-2\zeta_{2}\approx 1-{{12F_{,R}\left(\left[FF_{,RR}-F_{,R}^{2}\right]R+F_{,R}F\right)}\over{F_{,RR}\left(2F_{,R}R-F\right)^{2}}}~,\end{split}$ (82)

### v2

@@ 4.1 Einstein frame
$\varepsilon^{(E)}={}-{{1}\over{H_{E}^{2}}}{{dH_{E}}\over{dt_{E}}}$ (4.1)
@@
$\eta^{(E)}=\varepsilon^{(E)}-{{1}\over{2\,\varepsilon^{(E)}\,H_{E}}}{{d\varepsilon^{(E)}}\over{dt_{E}}}~.$ (4.2)
After the differentiation of Eq. (4.1) as
${{d\varepsilon^{(E)}}\over{dt_{E}}}=2H_{E}\,\left(\varepsilon^{(E)}\right)^{2}-{{1}\over{H^{2}_{E}}}{{d^{2}H_{E}}\over{dt_{E}^{2}}}~,$ (4.3)
@@
$\eta^{(E)}={}-{{1}\over{2\,H_{E}}}{{{{d^{2}H_{E}}\over{dt_{E}^{2}}}}\over{{{dH_{E}}\over{dt_{E}}}}}\,~.$ (4.4)
@@
${{d^{2}H_{E}}\over{dt_{E}^{2}}}={}-{{1}\over{M^{2}_{\mathrm{Pl}}}}{{d^{2}\phi}\over{dt_{E}^{2}}}\,{{d\phi}\over{dt_{E}}}\,$~, (4.5)
@@
$\eta^{(E)}={}-{{1}\over{H_{E}{{d\phi}\over{dt_{E}}}}}{{d^{2}\phi}\over{dt_{E}^{2}}}~.$ (4.6)
Next, substituting Eq. (4.6) into Eq. (3.20) yields the inflaton equation of motion in the form  [^2] Equation (4.7) in the case of the ultra-slow-roll (USR) regime with $V_{E,\phi}\longrightarrow 0$ implies $\eta^{(E)}\longrightarrow 3$. In the literature, the normalization of the SR parameter $\eta^{(E)}$ sometimes differs by the sign and the factor of 2.
$H_{E}{{d\phi}\over{dt_{E}}}\left(3-\eta^{(E)}\right)+V_{E,\phi}=0~.$ (4.7)
@@
in terms of e-folds $dN_{E}=-H_{E}dt=-dk/k$ or scale $k$. This is consistent with $N_{E}=-\ln(a_{E}/a_{E_{0}})$, where $a_{E}(t)$ is the FLRW scale factor in the Einstein frame and the constant $a_{E_{0}}$ is the value of $a_{E}$ at the end of inflation. Hence, $N_{E}$ is a monotonically decreasing function during inflation and $N_{E}=0$ corresponds to the end of inflation. The Planck pivot scale $k_{*}$ of the horizon crossing is taken at $k_{*}=0.05~{\rm Mpc}^{-1}$, which corresponds to $N_{E_{*}}=56\pm 8$ for Starobinsky inflation [41].
@@ 4.2 Jordan frame
$\varepsilon_{1}={}-{{\dot{H}}\over{H^{2}}}={{1}\over{2}}\,{{d\ln(H^{2})}\over{dN}}\,,\qquad\varepsilon_{n}={{\dot{\varepsilon}_{n-1}}\over{H{\varepsilon}_{n-1}}}={}-{{d\ln(\varepsilon_{n-1})}\over{dN}}~.$ (4.9)
We find it useful to introduce the additional SR parameters in our setup, which are given by
$\zeta_{1}={{\dot{F}_{,R}}\over{HF_{,R}}}={{F_{,RR}\dot{R}}\over{HF_{,R}}}={}-{{d\ln(F_{,R})}\over{dN}}\,,\quad\zeta_{n}={{\dot{\zeta}_{n-1}}\over{H{\zeta}_{n-1}}}={}-{{d\ln(\zeta_{n-1})}\over{dN}}\,.$ (4.10)
Using Eq. (2.8), the first SR parameters can be expressed in terms of $H^{2}$ and $R$ as
$\varepsilon_{1}=2-{{R}\over{6H^{2}}},\quad\zeta_{1}={{\left(6H^{2}-R\right)F_{,R}+F}\over{6H^{2}F_{,RR}\left(12H^{2}-R\right)}}-{{R}\over{3H^{2}}}~.$ (4.11)
Equation (2.9) can be rewritten to the form
$Y={{M_{\mathrm{Pl}}H}\over{\sqrt{2F_{,R}\,}}}\left(1+{{1}\over{2}}\zeta_{1}\right)~,$ (4.12)
and Eq. (2.11) can be rewritten to
$3M_{\mathrm{Pl}}^{4}H^{2}\left(1+{{1}\over{2}}\zeta_{1}\right)^{2}=AF_{,R}{\dot{R}}^{2}+2F_{,R}V_{\mathrm{eff}}~.$ (4.13)
@@
${{dH_{E}}\over{dt_{E}}}={{M_{\mathrm{Pl}}\dot{Y}}\over{\sqrt{2\,F_{,R}}}}={{M_{\mathrm{Pl}}^{2}H^{2}}\over{8F_{,R}}}\left(2\zeta_{1}\zeta_{2}-(2\varepsilon_{1}+\zeta_{1})(2+\zeta_{1})\right)\,$, (4.14)
@@
$\begin{split}\varepsilon^{(E)}&={}-{{1}\over{H_{E}^{2}}}\,{{d{H}_{E}}\over{d{t}_{E}}}=\varepsilon_{1}+{{\zeta_{1}\left(1-\varepsilon_{1}\right)}\over{2+\zeta_{1}}}-{{2\zeta_{1}\zeta_{2}}\over{\left(2+\zeta_{1}\right)^{2}}}\\ &\approx\varepsilon_{1}+{{1}\over{2}}\zeta_{1}-{{\zeta_{1}}\over{4}}\left(\zeta_{1}+2\zeta_{2}+2\varepsilon_{1}\right),\end{split}$ (4.15)
@@
Using Eq. (3.5) in the following form:
$\zeta_{1}={}-2\varepsilon_{1}+\zeta_{1}\left(\zeta_{1}+\zeta_{2}-\varepsilon_{1}\right)~,$ (4.16)
@@
$\varepsilon^{(E)}\approx\varepsilon_{1}+{{1}\over{2}}\zeta_{1}-{{\zeta_{1}}\over{4}}\left(\zeta_{1}+2\zeta_{2}+2\varepsilon_{1}\right)={{1}\over{4}}\zeta_{1}\left(\zeta_{1}-4\varepsilon_{1}\right)\approx 3\varepsilon_{1}^{2}+{\cal O}\left(\varepsilon_{1}^{3}\right).$ (4.17)
Similarly, by using Eqs. (4.2) and (4.17), we get
$\eta^{(E)}=\varepsilon^{(E)}+{{1}\over{2}}\,{{d\ln(\varepsilon^{(E)})}\over{dN_{E}}}\approx{}-\varepsilon_{2}{{dN}\over{dN_{E}}}+3\varepsilon_{1}^{2}+{\cal O}\left(\varepsilon_{1}^{3}\right)~.$ (4.18)
The relation
${{dN}\over{dN_{E}}}={{2}\over{2+\zeta_{1}}}\approx 1+\varepsilon_{1}$ (4.19)
gives
$\eta^{(E)}\approx{}-\varepsilon_{2}-\varepsilon_{1}\varepsilon_{2}+3\varepsilon_{1}^{2}+{\cal O}\left(\varepsilon_{1}^{3}\right)\,.$ (4.20)
@@
$r\!$ $\approx$ $\!16\varepsilon^{(E)}\approx 4\left|\zeta_{1}\left(\zeta_{1}-4\varepsilon_{1}\right)\right|\approx 48\varepsilon_{1}^{2}\,$, (4.21)
$n_{s}\!$ $\approx$ $\!1-4\varepsilon^{(E)}+2\eta^{(E)}\approx 1-2\varepsilon_{2}-6\varepsilon_{1}^{2}-2\varepsilon_{1}\varepsilon_{2}\approx 1-2\varepsilon_{2}\,$, (4.22)
$\alpha_{s}\!$ $\approx$ $\!{}-2\varepsilon_{2}\varepsilon_{3},\qquad\alpha_{t}\approx 2~{{d\,\varepsilon^{(E)}}\over{dN_{E}}}\approx{}-12\varepsilon^{2}_{1}\varepsilon_{2}.$ (4.23)
@@
$A_{s}={{H_{E}^{2}}\over{8\pi^{2}M_{\mathrm{Pl}}^{2}\varepsilon^{(E)}}}\approx{{H^{2}}\over{48\pi^{2}F_{,R}\varepsilon^{2}_{1}}}\approx{{H^{2}}\over{\pi^{2}rF_{,R}}}\,.$ (4.24)
Of course, Eqs. (4.21)–(4.24) make sense only in the SR approximation, because the higher-order terms with respect to $\varepsilon^{(E)}$ and $\eta^{(E)}$ are ignored.
@@ 4.3 USR regime
$A\ddot{\sigma}+{{A_{,\sigma}}\over{2}}{\dot{\sigma}}^{2}\ll V_{\mathrm{eff},\sigma}$ (4.25)
because the field derivative of the potential $V_{\mathrm{eff}}$ almost vanishes. Instead, Eq. (3.14) reads
$\ddot{\sigma}={}-3Y\sqrt{{{2F_{,R}}\over{M_{\mathrm{Pl}}^{2}}}}\dot{\sigma}-{{A_{,\sigma}}\over{2A}}{\dot{\sigma}}^{2}.$ (4.26)
The USR regime can be realized by engineering a nearly-inflection point in the inflaton potential below the inflation scale. It follows from Eq. (2.7) that at such (de Sitter-like) point one has
$H_{dS}^{2}={{F(R_{dS})}\over{6F_{,R}(R_{dS})}}\quad{\rm and,~hence,}\quad R_{dS}={{2F(R_{dS})}\over{F_{,R}(R_{dS})}}~.$ (4.27)
Under the condition (2.4), Eq. (4.27) is equivalent to
$V_{eff,\sigma}(\sigma_{dS})=0\quad{\rm and,~hence,}\quad F_{,\sigma}(\sigma_{dS})={{2F(\sigma_{dS})}\over{\sigma_{dS}}}\,.$ (4.28)
@@
$F_{,\sigma\sigma}(\sigma_{dS})={{2F(\sigma_{dS})}\over{\sigma_{dS}^{2}}}\,.$ (4.29)
@@
$V_{E}={{M^{2}_{\mathrm{Pl}}}\over{2}}R\exp\left(-\sqrt{{{2}\over{3}}}{{\phi}\over{M_{\mathrm{Pl}}}}\right)-F\exp\left(-2\sqrt{{{2}\over{3}}}{{\phi}\over{M_{\mathrm{Pl}}}}\right)~,$ (4.30)
$V_{E,\phi}=\sqrt{{{2}\over{3}}\,}{M_{\mathrm{Pl}}}^{-1}\left({F}\exp\left(-2\sqrt{{{2}\over{3}}}{{\phi}\over{M_{\mathrm{Pl}}}}\right)-V_{E}\right)~,$ (4.31)
@@
$V_{E,\phi\phi}=\sqrt{{{2}\over{3}}}\left[{{M_{\mathrm{Pl}}}\over{2}}\exp\left(-\sqrt{{{2}\over{3}}}{{\phi}\over{M_{\mathrm{Pl}}}}\right){{dR}\over{d\phi}}-{{2F}\over{M^{2}_{\mathrm{Pl}}}}\sqrt{{{2}\over{3}}}\exp\left(-2\sqrt{{{2}\over{3}}}{{\phi}\over{M_{\mathrm{Pl}}}}\right)-{{V_{E,\phi}}\over{M_{\mathrm{Pl}}}}\right]~.$ (4.32)
@@
$V_{E}\approx F\exp\left(-2\,\sqrt{{{2}\over{3}}}{{\phi}\over{M_{\mathrm{Pl}}}}\right)\approx{{F}\over{F_{,R}^{2}}}~.$ (4.33)
@@ 5 More on SR approximation in Jordan frame
Equation (2.11) in the SR approximation is just the Friedmann equation,
$3M_{\mathrm{Pl}}^{2}Y^{2}\approx V_{\mathrm{eff}}\,.$ (5.1)
After differentiating this equation with respect to the cosmic time in the Jordan frame and using Eq. (3.15), we obtain
$\dot{R}\approx{}-{{M_{\mathrm{Pl}}V_{\mathrm{eff},R}}\over{3YA\sqrt{2F_{,R}}}}\,~.$ (5.2)
@@
${R\,}^{\prime}\approx{{M_{\mathrm{Pl}}V_{\mathrm{eff},R}}\over{3YHA\sqrt{2F_{,R}}}}={{M_{\mathrm{Pl}}V_{\mathrm{eff},R}Y}\over{3Y^{2}HA\sqrt{2F_{,R}}}}\approx{{M_{\mathrm{Pl}}^{4}V_{\mathrm{eff},R}}\over{2F_{,R}AV_{\mathrm{eff}}}}\left(1-{{F_{,RR}}\over{2F_{,R}}}{R\,}^{\prime}\right)\,.$ (5.3)
After solving Eq. (5.3) for ${R\,}^{\prime}$, we get the SR evolution equation for the Ricci scalar $R$ in the Jordan frame as
${R\,}^{\prime}\approx{}-{{2F_{,R}\left(RF_{,R}-2F\right)}\over{F_{,RR}\left(2RF_{,R}-F\right)}}\,.$ (5.4)
@@
$\zeta_{1}(R)\approx{{2\left(RF_{,R}-2F\right)}\over{2RF_{,R}-F}}$ (5.5)
@@
$H^{2}(R)\approx{{\left(2RF_{,R}-F\right)^{2}}\over{54F_{R}\left(RF_{R}-F\right)}}~.$ (5.6)
Equation (4.11) also implies
$\varepsilon_{1}=2-{{R}\over{6H^{2}}}\approx 2-{{9F_{R}R\left(RF_{R}-F\right)}\over{\left(2RF_{,R}-F\right)^{2}}}\,~.$ (5.7)
Using Eqs. (5.5) and (5.7), we find in the SR approximation that
$\varepsilon_{1}(R)={}-{{1}\over{2}}\zeta_{1}(R)+{{1}\over{4}}\zeta_{1}^{2}(R).$ (5.8)
Then, by using Eqs. (5.4), (5.5) and (5.7), we get
$\varepsilon_{2}(R)={}-{{1}\over{\varepsilon_{1}}}{{d\varepsilon_{1}}\over{dN}}={}-{{{R\,}^{\prime}}\over{\varepsilon_{1}}}{{d\varepsilon_{1}}\over{dR}}\\ \approx{{18FF_{,R}\left(FF_{,RR}R-F_{,R}^{2}R+F_{,R}F\right)}\over{F_{,RR}\left(4F_{,R}^{3}R^{3}-3F^{2}F_{,R}R+F^{3}\right)}}~,$ (5.9)
$\zeta_{2}(R)\approx{{6F_{,R}\left(\left[FF_{,RR}-F_{,R}^{2}\right]R+F_{,R}F\right)}\over{F_{,RR}\left(2F_{,R}R-F\right)^{2}}}~,$ (5.10)
@@
$\begin{split}\zeta_{3}(R)&\approx{}-{{4\left(F_{,R}R-2\,F\right)}\over{F_{,RR}^{2}\left(F_{,RR}FR-F_{,R}^{2}R+F_{,R}F\right)\left(2F_{,R}R-F\right)^{2}}}\\ &\times\left(F_{,RR}\left[{{F}\over{2}}F_{,RR}^{2}R\left(2F_{,R}R+F\right)-{{F}\over{2}}F_{,R}F_{,RR}\left(2F_{,R}R-3F\right)\right.\right.\\ &\left.\left.{}+F_{,R}^{3}\left(F-F_{,R}R\right)\right]-{{1}\over{2}}F_{,RRR}F_{,R}^{2}\left(F_{,R}R-F\right)\left(2F_{,R}R-F\right)\right)\,.\end{split}$ (5.11)
Equation (5.8) in the same approximation implies
${{d\epsilon_{1}}\over{dN}}\approx-{{1}\over{2}}{{d\zeta_{1}}\over{dN}}+{{1}\over{4}}{{d\zeta_{1}^{2}}\over{dN}}\approx{{\zeta_{1}\zeta_{2}}\over{2}}-{{\zeta_{1}^{2}\zeta_{2}}\over{2}}\approx{{\zeta_{1}}\over{2}}\left(\zeta_{2}-\zeta_{1}\zeta_{2}\right)~.$ (5.12)
@@
$\epsilon_{2}\approx\zeta_{2}(1-\zeta_{1})\approx\zeta_{2}$ (5.13)
@@
$\epsilon_{3}\approx\zeta_{3}-{{\zeta_{1}\zeta_{2}}\over{1-\zeta_{1}}}\approx\zeta_{3}.$ (5.14)
We use the simplified relations $\varepsilon_{1}\approx{}-\zeta_{1}/2$, $\varepsilon_{2}\approx\zeta_{2}$, and $\varepsilon_{3}\approx\zeta_{3}$ to calculate the inflationary parameters by Eqs. (4.21)–(4.23), and Eq. (5.7) to get the value of $R$ corresponding to the end of inflation. Substituting those results into the right-hand-side of Eq. (4.21) gives
$r\approx 48\varepsilon_{1}^{2}\approx 12\zeta_{1}^{2}\approx 48\left({{RF_{,R}-2F}\over{2RF_{,R}-F}}\right)^{2}$ (5.15)
@@
$\begin{split}n_{s}\approx 1-2\zeta_{2}\approx 1-{{12F_{,R}\left(\left[FF_{,RR}-F_{,R}^{2}\right]R+F_{,R}F\right)}\over{F_{,RR}\left(2F_{,R}R-F\right)^{2}}}~,\end{split}$ (5.16)